OpenStudy (smrougraff):

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B has a volume of 3456 cubic inches. How many times bigger is the volume of pyramid B than pyramid A? Give your answer as a percent. Provide an explanation and proof for your answer to receive full credit.

OpenStudy (smrougraff):

I know the answer to this, I just am not sure how to show it as a percentage.

OpenStudy (raden):

find the volume of pyramid A, first. Volume = 1/3 * base area * height = 1/3 * 12 * 12 * 8 = 384 inc^3

OpenStudy (smrougraff):

Ok I'm right so far

OpenStudy (raden):

then divided the biggest to smalest. VB/VA = 3456/384 = 9

OpenStudy (smrougraff):

So, that would mean that Pyramid B is 9 times bigger than Pyramid A. What about the percent?

OpenStudy (raden):

yes, correct. then just times by 100% 9 * 100% = 900%

OpenStudy (smrougraff):

Oh, I see. That's a lot easier than I thought! Thanks so much for your help!

OpenStudy (raden):

you're welcome